Circular No. 8665

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 B1 (McNAUGHT) R. H. McNaught reports the discovery of a slightly diffuse comet (dis-

covery observation tabulated below) on Siding Spring Survey images taken with the 0.5-m Uppsala Schmidt telescope. Some images from Jan. 27.6 UT suggest that the coma, which has a diameter of ∼ 8′′, extends to the west; McNaught adds that 60-s Uppsala Schmidt exposures taken by G. Garradd (possibly affected by light cirrus clouds) on Jan. 28.7 also show the object to be rather more diffuse than stars of similar brightness, with a faint coma of diameter ∼ 15′′ possibly extended towards the west. Three co-added 200-s CCD exposures in poor seeing by A. C. Gilmore (Mt. John 0.6-m f/7.6 reflector) on Jan. 30.45 and 30.46 show an uncondensed circular coma ∼ 13′′ diameter, with no hint of tail.

2006 UT  α_{2000}  δ_{2000}  Mag.
Jan. 27.62223  13^{h}08^{m}10.^{s}75  −39^{o}23′36′′.2  18.3

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2006-B86.

T = 2005 Nov. 3.589 TT  ω = 298.^{o}193
Ω = 161.411 }2000.0
q = 2.96619 AU  i = 134.175

SUPERNOVA 1985U S. V. Antipin, Sternberg Astronomical Institute and Institute of As-

tronomy (Russian Academy of Sciences), reports the discovery of an apparent supernova on archival plates in the Moscow collection that were taken with the Crimean 40-cm astrograph. The coordinates of the apparent supernova, measured relative to USNO-A2.0 stars, are α = 2^{h}25^{m}41^{s}.95, δ = +39^{o}35′12′′.3 (equinox 2000.0), which is 1′′ north of the galaxy’s position in the USNO-A2.0 catalogue. The star is seen on five plates, at the following photographic magnitudes: 1985 Jan. 13.75 UT, [16.2; Sept. 16.03, 15.2; 17.02, 15.0; 21.99, 15.2; 22.02, 15.3; 25.03, 15.9:; 1986 Sept. 3.98, [17.0. SN 1985U is not present on Digitized Sky Survey images or on any other plates of the Moscow collection. The parent galaxy is difficult to see on the best Moscow plates as a nebular object at m_{pg} ∼ 17.0.

2006 January 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green